OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -9 - 4 + 1 + 6 + ... + 66

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@Reaper534

OpenStudy (anonymous):

ohh thats easy it adds by 5

OpenStudy (anonymous):

so 11 16 21 26 31 ect.

OpenStudy (anonymous):

so is it -9+5n

OpenStudy (anonymous):

\[\sum_{n=0}^{15} (-9+5n)\]

OpenStudy (anonymous):

@Reaper534

OpenStudy (anonymous):

tht will be the best choice

OpenStudy (anonymous):

thank you can you also help me in this problem

OpenStudy (anonymous):

Find the sum of the first 9 terms of the sequence. Show all work for full credit. 2, -7, -16, -25, ...

OpenStudy (anonymous):

2-9=-7-8=-16-... so u subtract by 9 im alo in school right now too i need to get back at that lol

OpenStudy (anonymous):

lol thank you so i just subtract -16-9

OpenStudy (anonymous):

yeah u do u start from 2 or 25?

OpenStudy (anonymous):

alright so 2-9, -7-9, -16-9, 25-9 and i will have my answer

OpenStudy (anonymous):

2-9=7,...

OpenStudy (anonymous):

ok gtg to my paper

OpenStudy (anonymous):

-7-9= -16, -16-9=-25-9=-34... and so on

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

an i do this unitl i get to the ninth term

OpenStudy (anonymous):

ok there u go now i relly gtg

OpenStudy (anonymous):

ight thanks